# RxnBench: A Multimodal Benchmark for Evaluating Large Language Models on Chemical Reaction Understanding from Scientific Literature

### v2

@@ Introduction
The past decade has witnessed a paradigm shift in scientific discovery driven by Artificial Intelligence (AI), particularly in chemistry, the central science, where data-driven approaches are reshaping traditional research workflows Wang2023. From early Quantitative Structure-Activity Relationship (QSAR) models Cherkasov2014 to modern deep learning-based molecular generation GomezBombarelli2018 and property prediction Yang2019, AI has demonstrated potential that often transcends human intuition. Notably, the breakthrough in protein structure prediction achieved by geometric deep learning models like AlphaFold Jumper2021 has significantly inspired the chemical community to explore AI applications in small molecule drug discovery Vamathevan2019, materials design Merchant2023, and complex organic retrosynthesis planning Segler2018; Coley2019.
@@
In recent years, the emergence of Large Language Models (LLMs) has introduced a new variable to chemical research. Unlike specialized models focused on single tasks, LLMs exhibit surprising general reasoning capabilities and knowledge integration skills OpenAI2023. Through instruction tuning or agent-based frameworks (such as ChemCrow Bran2023), LLMs have been proven capable of performing literature retrieval, experimental planning, and even controlling automated laboratory equipment Boiko2023. However, most existing chemical large models rely heavily on linearized text representations (such as SMILES Weininger1988 or SELFIES Krenn2022) to process chemical information. Although the textual modality is a significant carrier of knowledge, the core knowledge in chemical literature is often highly condensed within the unstructured visual modality, limiting the complete acquisition and understanding of chemical knowledge by text-only LLMs.
@@
There are unique challenges in chemical literature understanding. Achieving machine reading of chemical literature faces multiple challenges not encountered in general document understanding tasks: First, the chemical reaction scheme is a highly condensed domain-specific graphical language. Unlike natural images, a reaction scheme is a complex semantic network composed of molecular structures, arrow symbols, reaction conditions (temperature, solvents, catalysts, etc.), and text labels. As shown in Figure 1a, understanding a reaction scheme requires not only recognizing discrete molecular structures but also inferring their roles in a dynamic chemical process (reactants, products, intermediates, or transition states). This type of “process understanding” is far more complex than simple object detection. Second, the unstructured nature of PDF documents poses significant engineering and cognitive hurdles. Scientific literature is primarily disseminated in PDF format—a format designed for human visual consumption rather than machine parsing—often referred to as a “graveyard” for data Tkaczyk2015. In chemical papers, reaction scheme images are frequently spatially separated from their textual descriptions, requiring models to handle cross-page layout dependencies. Third, the strong association and complementarity of multimodal information (Strong Multimodality). Chemical knowledge rarely exists in a single modality in isolation. For instance, a code in the main text (e.g., “Compound 3a”) must be aligned with the structure in an illustration (e.g., “Scheme 1”), while specific yields or enantiomeric excess ($e$) values might be hidden in a supplementary table (e.g., “Table S2”). Models require the ability to perform joint reasoning across text, images, and tables Yin2024. Finally, the extreme requirement for precision. Chemistry is an exact science. In natural image captioning, misidentifying a “dog” as a “cat” is an error; however, in chemistry, a mere change in the direction of a chemical bond (wedge vs. dash) implies a reversal of stereochemistry, which in medicinal chemistry can mean the difference between a cure and a poison (e.g., the thalidomide tragedy). Existing multimodal models suffer from severe “hallucination” issues Nori2023, tending to generate structures that appear plausible but are chemically incorrect.
@@
As for the existing benchmarks for chemistry, we categorize them into two types based on input modality: single-modality benchmarks, which present problems purely in textual form, and multimodality benchmarks, which incorporate both textual and visual inputs, such as molecular structures or crystallographic images. Representative single-modal benchmarks include ChemBench Mirza2024, ChemLLMBench Guo2023, MMLU Hendrycks2020, SciBench Wang2023SciBench, SciCode Tian2024, JEEBench Arora2023, and OlympicArena He2024. While these provide robust evaluation for textual reasoning, they inherently lack the ability to assess visual chemical literacy. Conversely, multimodal initiatives such as ScienceQA Saikh2022, OlympiadBench He2024, ChemTable Zhou2025, Mol-Puzzle Guo2024, and ScemQA Liang2024 have integrated visual data to assess broader scientific literacy. Yet, a critical gap remains: these benchmarks primarily target static concept recognition or elementary interpretation, often treating molecular graphs as isolated objects rather than components of a dynamic reaction process. They generally lack the granularity required to evaluate the precise structural and quantitative reasoning—such as evaluating reaction yields of different products, or verifying stereochemical fidelity—that is central to organic synthesis. Furthermore, none of these datasets challenge models to synthesize information from the full, unstructured context of reaction-containing PDF literatures, a capability that fundamentally distinguishes our work from existing concept-centric evaluations. This limitation motivates the development of RxnBench, which aims to address these gaps through broader coverage of reaction types and a distinct focus on reaction perception and full-document reasoning.
@@
Given the subtle nuances of organic chemistry—where a single stereocenter inversion or a misinterpreted abbreviation can invalidate a reaction—crowdsourcing platforms were deemed unsuitable for this task. Instead, we recruited a team of domain experts to perform the annotation. The annotation team consisted of doctoral candidates (Ph.D. students) specializing in organic chemistry. All annotators possessed extensive experience in reading synthetic literature and were proficient in using chemical drawing software and interpreting mechanism of organic reactions. Furthermore, they were skilled in Optical Chemical Structure Recognition (OCSR) workflows, to convert graphical structures directly into Extended SMILES (E-SMILES) Fang2024, facilitating precise structural editing and verification.
@@
Figure Extraction and Segmentation is the first step. We utilized Uni-Parser Fang2025, a specialized layout analysis tool for scientific documents, to process the source PDFs. Reaction schemes were automatically detected, cropped, and saved as high-resolution single images. Crucially, this process was designed to identify and include the scheme's associated titles and labels (subscripts) within the cropped area, thereby ensuring the completeness of the visual information. This approach guarantees that the visual inputs for the benchmark contain all necessary internal context while being isolated from the surrounding main text, focusing the task strictly on visual chemical perception and preventing information leakage from the body text.
The second step involves Hybrid Question Generation, where we adopted a “Human-in-the-Loop” strategy to produce high-quality Question-Answer (QA) pairs. Initially, OpenAI o4-mini Hurst2024 was seeded with specific chemical templates to propose a set of five candidate QA pairs per image. (The prompt was documented in Supporting Information S4). Domain experts then acted as “exam setters”, rigorously reviewing, modifying, or completely rewriting these questions to ensure chemical relevance. These questions were strictly categorized into two tiers to assess different levels of cognitive capability. The first tier, Descriptive Questions, focuses on information extraction, assessing the model's ability to accurately “read” the diagram. This encompasses tasks such as identifying reaction titles, enumerating reactants and products, and extracting precise reaction parameters like temperature, reaction time, and specific activation methods (e.g., microwave irradiation, ball milling, or photoredox conditions). It also involves distinguishing the roles of various species, such as differentiating a metal catalyst from its ligand. The second tier, Reasoning Questions, requires synthesizing visual information with chemical knowledge to derive conclusions. This ranges from simple reasoning, such as comparing yields and enantioselectivity across substrates or performing Markush enumeration, to complex reasoning tasks like mechanism analysis, identifying bond cleavage sites, and tracing catalytic cycles. Advanced inquiries further test the model's ability to analyze structure-activity relationships (SAR) based on electronic effects and perform retrosynthetic planning.
@@ Acknowledgements
Table S1: Performance comparison of representative MLLMs on the SF-QA.
Model Think Weight RxnBench-En RxnBench-Zh Mean Score
Gemini-3-Flash-preview - 0.9593 0.9652 0.9623
Seed1.8-Think - 0.9325 0.9403 0.9364
Gemini-3-Pro-preview - 0.9318 0.9403 0.9361
GPT-5(high) - 0.9279 0.9246 0.9263
Gemini-2.5-Pro - 0.9095 0.9423 0.9259
GPT-5.1(high) - 0.9213 0.9220 0.9216
GPT-5(medium) - 0.9207 0.9226 0.9216
Qwen3-VL-235B-A22B-Think Open 0.9220 0.9134 0.9177
Qwen3-VL-32B-Think Open 0.9128 0.9161 0.9144
GPT-5.1(medium) - 0.9108 0.9141 0.9125
GPT-5-mini - 0.9108 0.9128 0.9118
Seed1.5-VL-Think - 0.9056 0.9161 0.9109
o3 - 0.9056 0.9115 0.9086
o4-mini - 0.9062 0.9075 0.9069
InternVL3.5-241B-A28B Open 0.9003 0.9062 0.9033
Intern-S1 Open 0.8938 0.8944 0.8941
Qwen3-VL-30B-A3B-Think Open 0.8689 0.8590 0.8689
Qwen3-VL-Plus - 0.8551 0.8656 0.8604
Qwen3-VL-8B-Think Open 0.8636 0.8564 0.8600
Seed1.5-VL - 0.8518 0.8669 0.8594
Qwen3-VL-235B-A22B-Instruct Open 0.8492 0.8675 0.8584
InternVL3-78b Open 0.8531 0.8308 0.8420
Qwen3-VL-4B-Think Open 0.8577 0.8256 0.8416
Intern-S1-mini Open 0.8521 0.8282 0.8402
GLM-4.1V-9B-Thinking Open 0.8392 0.8341 0.8367
Qwen3-VL-32B-Instruct Open 0.8315 0.8407 0.8361
Qwen2.5-VL-72B Open 0.8341 0.8308 0.8325
Qwen2.5-VL-Max - 0.8192 0.8262 0.8227
GPT-5-nano - 0.7980 0.7941 0.7961
Qwen2.5-VL-32B Open 0.7980 0.7908 0.7944
Gemini-2.5-Flash - 0.6925 0.8557 0.7741
Qwen3-VL-8B-Instruct Open 0.7548 0.7495 0.7521
Qwen3-VL-30B-A3B-Instruct Open 0.7456 0.7436 0.7456
GPT-4o - 0.7462 0.7436 0.7449
Qwen2.5-VL-7B Open 0.7082 0.7233 0.7158
Qwen3-VL-4B-Instruct Open 0.7023 0.7023 0.7023
Qwen3-VL-2B-Think Open 0.6780 0.6708 0.6744
Qwen2.5-VL-3B Open 0.6748 0.6643 0.6696
GPT-4o mini - 0.6636 0.6066 0.6351
Qwen3-VL-2B-Instruct Open 0.5711 0.5928 0.5820
Choose the longest answer - - 0.4262 0.4525 0.4394
Deepseek-VL2 Open 0.4426 0.4216 0.4321
Random - - 0.2500 0.2500 0.2500
Table S2: Performance of evaluated models across SF-QA Task Types
Model Thinka Weightb F.E.c RR&F.d M&P.e CA&R.f G.U.g S.R.h
Gemini-3-Flash-preview - 0.9613 0.9643 0.9764 0.9630 0.9492 0.9030
Seed1.8-Think - 0.9331 0.9484 0.9527 0.9444 0.9492 0.8284
Gemini-3-Pro-preview - 0.9648 0.9246 0.9527 0.9398 0.9322 0.7463
GPT-5(high) - 0.9313 0.9444 0.9527 0.9167 0.9661 0.8358
Gemini-2.5-Pro - 0.9331 0.9246 0.9459 0.9491 0.9322 0.6343
GPT-5.1(high) - 0.9243 0.9524 0.9426 0.9167 0.9661 0.7910
GPT-5(medium) - 0.9349 0.9325 0.9493 0.9167 0.9492 0.7761
Qwen3-VL-32B-Think Open 0.9296 0.9405 0.9426 0.9259 0.9153 0.7015
GPT-5.1(medium) - 0.9243 0.9365 0.9426 0.9167 0.9492 0.7090
GPT-5-mini - 0.9225 0.9325 0.9257 0.9259 0.9831 0.7388
Seed1.5-VL-Think - 0.8996 0.9365 0.9358 0.9074 0.9153 0.8060
o3 - 0.9313 0.9325 0.9223 0.8981 0.9492 0.7090
o4-mini - 0.6391 0.7302 0.7500 0.6667 0.6271 0.4627
Intern-S1 Open 0.9014 0.9127 0.9223 0.9028 0.8814 0.7463
Qwen3-VL-30B-A3B-Think Open 0.8732 0.8810 0.9054 0.8843 0.9322 0.6940
Qwen3-VL-Plus - 0.8275 0.8968 0.8986 0.8565 0.9153 0.7687
Qwen3-VL-8B-Think Open 0.8768 0.8730 0.8885 0.9028 0.8983 0.6567
Seed1.5-VL - 0.9327 0.9127 0.9122 0.8472 0.8305 0.7015
InternVL3-78b Open 0.8556 0.8730 0.8885 0.8981 0.9153 0.6194
Qwen3-VL-4B-Think Open 0.8838 0.8770 0.8615 0.9074 0.8983 0.6045
Intern-S1-mini Open 0.8239 0.8690 0.8547 0.8611 0.8475 0.6791
GLM-4.1V-9B-Thinking Open 0.8433 0.8690 0.8649 0.8657 0.8814 0.6493
Qwen3-VL-32B-Instruct Open 0.8169 0.8571 0.8885 0.8519 0.8305 0.6866
Qwen2.5-VL-72B Open 0.8063 0.8063 0.8770 0.9088 0.8102 0.9322
Qwen2.5-VL-Max - 0.7958 0.8571 0.8885 0.8194 0.8983 0.6642
GPT-5-nano - 0.8063 0.8452 0.8311 0.8241 0.7797 0.5672
Qwen2.5-VL-32B Open 0.7729 0.8413 0.8750 0.8009 0.8305 0.6418
Gemini-2.5-Flash - 0.7799 0.6111 0.6757 0.6620 0.7627 0.5373
Qwen3-VL-8B-Instruct Open 0.7113 0.8175 0.8446 0.8241 0.7627 0.5075
Qwen3-VL-30B-A3B-Instruct Open 0.7042 0.7937 0.8311 0.7824 0.7119 0.5970
GPT-4o - 0.7359 0.8175 0.7973 0.7500 0.7627 0.5224
Qwen2.5-VL-7B Open 0.6678 0.7659 0.8041 0.7130 0.6441 0.5373
Qwen3-VL-4B-Instruct Open 0.6708 0.7302 0.7804 0.7222 0.6610 0.5970
Qwen3-VL-2B-Think Open 0.7342 0.6706 0.7128 0.7083 0.6102 0.3657
Qwen2.5-VL-3B Open 0.6426 0.7381 0.7635 0.6898 0.6610 0.4776
GPT-4o mini - 0.6391 0.7302 0.7500 0.6667 0.6271 0.4627
Qwen3-VL-2B-Instruct Open 0.5405 0.6190 0.6318 0.6250 0.6102 0.3731
Deepseek-VL2 Open 0.4120 0.5040 0.4899 0.4907 0.3729 0.3060
Random - - 0.2500 0.2500 0.2500 0.2500 0.2500 0.2500
  • a
  • b
  • c
    F.E.: Fact Extraction
  • d
    RR&F.: Reagent Roles & Functions
  • e
    M&P.: Mechanism & Process
  • f
    CA&R.: Comparative Analysis & Reasoning
  • g
    G.U.: Global Understanding
  • h
    S.R.: Structure Recognition
Table S3: Performance of MLLMs across FD-QA Task Types
Model Thinka Weightb Rxn.En.c Rxn.Zh.d C.R.e S.R.f O.S.g
Gemini-3-Flash-preview - 0.4574 0.4685 0.5530 0.4264 0.4630
Gemini-3-Pro-preview - 0.4167 0.4407 0.5000 0.3897 0.4287
Gemini-2.5-Pro - 0.4148 0.4426 0.5682 0.3652 0.4287
Gemini-2.5-Flash - 0.3519 0.3815 0.4394 0.3235 0.3667
Seed1.8-Think - 0.3352 0.3889 0.4242 0.3064 0.3620
GPT-5(high) - 0.3444 0.3463 0.4470 0.3113 0.3454
GPT-5.1(high) - 0.3333 0.3426 0.4545 0.2941 0.3380
Seed1.5-VL-Think - 0.3148 0.3574 0.3939 0.2892 0.3361
Qwen3-VL-32B-Think Open 0.2981 0.3167 0.4015 0.2647 0.3074
Qwen3-VL-235B-A22B-Think Open 0.3056 0.3056 0.4091 0.2721 0.3056
Qwen3-VL-30B-A3B-Think Open 0.2611 0.2481 0.3409 0.2353 0.2546
Qwen3-VL-8B-Think Open 0.2333 0.2315 0.3333 0.2010 0.2324
Qwen3-VL-4B-Think Open 0.2352 0.1870 0.3106 0.2108 0.2111
Intern-S1 Open 0.1963 0.1944 0.2652 0.1740 0.1954
Seed1.5-VL-Instruct x - 0.1630 0.1741 0.2045 0.1495 0.1685
Qwen3-VL-2B-Think Open 0.1444 0.1463 0.1818 0.1324 0.1453
GPT-4o - 0.1426 0.1389 0.1515 0.1397 0.1407
Qwen3-VL-32B-Instruct Open 0.1241 0.0981 0.1894 0.1029 0.1111
Qwen3-VL-8B-Instruct Open 0.0907 0.1019 0.1591 0.0686 0.0963
Qwen3-VL-30B-A3B-Instruct Open 0.0778 0.0815 0.1288 0.0613 0.0796
Qwen3-VL-4B-Instruct Open 0.0611 0.0611 0.0606 0.0612 0.0611
Qwen3-VL-2B-Instruct Open 0.0481 0.0667 0.0758 0.0392 0.0574
  • a
  • b
  • c
    Rxn.En.: RxnBench-Doc-En
  • d
    Rxn.Zh.: RxnBench-Doc-Zh
  • e
    C.R.: Context Reasoning (En)
  • f
    S.R.: Structure Reasoning (En)
  • g
    O.S.: Overall Score

### v3

@@ Introduction
The past decade has witnessed a paradigm shift in scientific discovery driven by Artificial Intelligence (AI), particularly in chemistry, the central science, where data-driven approaches are reshaping traditional research workflows Wang et al. (2023a). From early Quantitative Structure-Activity Relationship (QSAR) models Cherkasov et al. (2014) to modern deep learning-based molecular generation Gómez-Bombarelli et al. (2018) and property prediction Yang et al. (2019), AI has demonstrated potential that often transcends human intuition. Notably, the breakthrough in protein structure prediction achieved by geometric deep learning models like AlphaFold Jumper et al. (2021) has significantly inspired the chemical community to explore AI applications in small molecule drug discovery Vamathevan et al. (2019), materials design Merchant et al. (2023), and complex organic retrosynthesis planning Segler et al. (2018); Coley et al. (2019).
@@
In recent years, the emergence of Large Language Models (LLMs) has introduced a new variable to chemical research. Unlike specialized models focused on single tasks, LLMs exhibit surprising general reasoning capabilities and knowledge integration skills Achiam et al. (2023). Through instruction tuning or agent-based frameworks (such as ChemCrow Bran et al. (2023)), LLMs have been proven capable of performing literature retrieval, experimental planning, and even controlling automated laboratory equipment Boiko et al. (2023). However, most existing chemical large models rely heavily on linearized text representations (such as SMILES Weininger (1988) or SELFIES Krenn et al. (2022)) to process chemical information. Although the textual modality is a significant carrier of knowledge, the core knowledge in chemical literature is often highly condensed within the unstructured visual modality, limiting the complete acquisition and understanding of chemical knowledge by text-only LLMs.
@@
There are unique challenges in chemical literature understanding. Achieving machine reading of chemical literature faces multiple challenges not encountered in general document understanding tasks: First, the chemical reaction scheme is a highly condensed domain-specific graphical language. Unlike natural images, a reaction scheme is a complex semantic network composed of molecular structures, arrow symbols, reaction conditions (temperature, solvents, catalysts, etc.), and text labels. As shown in Figure 1a, understanding a reaction scheme requires not only recognizing discrete molecular structures but also inferring their roles in a dynamic chemical process (reactants, products, intermediates, or transition states). This type of “process understanding” is far more complex than simple object detection. Second, the unstructured nature of PDF documents poses significant engineering and cognitive hurdles. Scientific literature is primarily disseminated in PDF format—a format designed for human visual consumption rather than machine parsing—often referred to as a “graveyard” for data Tkaczyk et al. (2015). In chemical papers, reaction scheme images are frequently spatially separated from their textual descriptions, requiring models to handle cross-page layout dependencies. Third, the strong association and complementarity of multimodal information (Strong Multimodality). Chemical knowledge rarely exists in a single modality in isolation. For instance, a code in the main text (e.g., “Compound 3a”) must be aligned with the structure in an illustration (e.g., “Scheme 1”), while specific yields or enantiomeric excess ($e$) values might be hidden in a supplementary table (e.g., “Table S2”). Models require the ability to perform joint reasoning across text, images, and tables Yin et al. (2024). Finally, the extreme requirement for precision. Chemistry is an exact science. In natural image captioning, misidentifying a “dog” as a “cat” is an error; however, in chemistry, a mere change in the direction of a chemical bond (wedge vs. dash) implies a reversal of stereochemistry, which in medicinal chemistry can mean the difference between a cure and a poison (e.g., the thalidomide tragedy). Existing multimodal models suffer from severe “hallucination” issues Nori et al. (2023), tending to generate structures that appear plausible but are chemically incorrect.
@@
As for the existing benchmarks for chemistry, we categorize them into two types based on input modality: single-modality benchmarks, which present problems purely in textual form, and multimodality benchmarks, which incorporate both textual and visual inputs, such as molecular structures or crystallographic images. Representative single-modal benchmarks include ChemBench Mirza et al. (2024), ChemLLMBench Guo et al. (2023), MMLU Hendrycks et al. (2020), SciBench Wang et al. (2023b), SciCode Tian et al. (2024), JEEBench Arora et al. (2023), and OlympicArena He et al. (2024). While these provide robust evaluation for textual reasoning, they inherently lack the ability to assess visual chemical literacy. Conversely, multimodal initiatives such as ScienceQA Saikh et al. (2022), OlympiadBench He et al. (2024), ChemTable Zhou et al. (2025), Mol-Puzzle Guo et al. (2024), and ScemQA Liang et al. (2024) have integrated visual data to assess broader scientific literacy. Yet, a critical gap remains: these benchmarks primarily target static concept recognition or elementary interpretation, often treating molecular graphs as isolated objects rather than components of a dynamic reaction process. They generally lack the granularity required to evaluate the precise structural and quantitative reasoning—such as evaluating reaction yields of different products, or verifying stereochemical fidelity—that is central to organic synthesis. Furthermore, none of these datasets challenge models to synthesize information from the full, unstructured context of reaction-containing PDF literatures, a capability that fundamentally distinguishes our work from existing concept-centric evaluations. This limitation motivates the development of RxnBench, which aims to address these gaps through broader coverage of reaction types and a distinct focus on reaction perception and full-document reasoning.
@@
Given the subtle nuances of organic chemistry—where a single stereocenter inversion or a misinterpreted abbreviation can invalidate a reaction—crowdsourcing platforms were deemed unsuitable for this task. Instead, we recruited a team of domain experts to perform the annotation. The annotation team consisted of doctoral candidates (Ph.D. students) specializing in organic chemistry. All annotators possessed extensive experience in reading synthetic literature and were proficient in using chemical drawing software and interpreting mechanism of organic reactions. Furthermore, they were skilled in Optical Chemical Structure Recognition (OCSR) workflows, to convert graphical structures directly into Extended SMILES (E-SMILES) Fang et al. (2025b), facilitating precise structural editing and verification.
@@
Figure Extraction and Segmentation is the first step. We utilized Uni-Parser Fang et al. (2025a), a specialized layout analysis tool for scientific documents, to process the source PDFs. Reaction schemes were automatically detected, cropped, and saved as high-resolution single images. Crucially, this process was designed to identify and include the scheme's associated titles and labels (subscripts) within the cropped area, thereby ensuring the completeness of the visual information. This approach guarantees that the visual inputs for the benchmark contain all necessary internal context while being isolated from the surrounding main text, focusing the task strictly on visual chemical perception and preventing information leakage from the body text.
The second step involves Hybrid Question Generation, where we adopted a “Human-in-the-Loop” strategy to produce high-quality Question-Answer (QA) pairs. Initially, OpenAI o4-mini Hurst et al. (2024) was seeded with specific chemical templates to propose a set of five candidate QA pairs per image. (The prompt was documented in Supporting Information S4). Domain experts then acted as “exam setters”, rigorously reviewing, modifying, or completely rewriting these questions to ensure chemical relevance. These questions were strictly categorized into two tiers to assess different levels of cognitive capability. The first tier, Descriptive Questions, focuses on information extraction, assessing the model's ability to accurately “read” the diagram. This encompasses tasks such as identifying reaction titles, enumerating reactants and products, and extracting precise reaction parameters like temperature, reaction time, and specific activation methods (e.g., microwave irradiation, ball milling, or photoredox conditions). It also involves distinguishing the roles of various species, such as differentiating a metal catalyst from its ligand. The second tier, Reasoning Questions, requires synthesizing visual information with chemical knowledge to derive conclusions. This ranges from simple reasoning, such as comparing yields and enantioselectivity across substrates or performing Markush enumeration, to complex reasoning tasks like mechanism analysis, identifying bond cleavage sites, and tracing catalytic cycles. Advanced inquiries further test the model's ability to analyze structure-activity relationships (SAR) based on electronic effects and perform retrosynthetic planning.